Diffusion Morning Report 02/02/05

Ted Berg tberg at snf.stanford.edu
Wed Feb 2 06:30:36 PST 2005


Good Morning All,
       Looking good this morning.  Implanter is out of Arsine we should 
get it this morning. Testing Teos2.  MFC ordered for SiGe,
Ebeam still in high vac.  Everything else looks OK for now,

. 	Date: 	2/2/2005 	 
Tube Name/Number 	Status 	Comments 	THICKNESS-Change at
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	down 	Qual testing 	 
Tylan 8-SiGe 	yellow 	SiH4 leak 	12.33.0/30
Tylan 9-Poly 	Up 	  	17.69/30
Tylan 10-Nitride 	Up 	  	11.1/30
Tylan 11-BPSG-LTO 	Up 	  	17.69/30
Tylan 12-FGA 	Up 	  	 
Tystar 	yellow 	  	3.48/30
Doped Poly 	GONE 	  	 
TEOS 	up 	  	 
Tylan 16 	N/A 	  	 
Thermco 1 	Up 	  	 
Thermco 2 	Up 	  	 
Thermco Poly 	yellow 	testing doped 	 
Thermco Nitride 	Down 	Waiting 	 
  	  	  	 
Burn Box 	Up 	  	 
ZYGO 	up 	  	 
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	Auto changeover working well 	A on line 480 PSI


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20050202/38fee418/attachment.html>


More information about the diffusion mailing list